OpenStudy (anonymous):

3 ln x-5 ln c ( Single Natural Logarithm)

OpenStudy (anonymous):

is that \[\huge 3\ln(x)-5\ln(c)\]?

OpenStudy (anonymous):

guess i am asking is it both a "c" and an "x" in there ?

OpenStudy (here_to_help15):

a*ln(b) = ln(b^a) and ln(b) + ln(c) = ln(bc) so 3ln(5) + ln(x) = ln(5^3) + ln(x) = ln(125x)

OpenStudy (anonymous):

Thank you, @Here_to_Help15 !

OpenStudy (anonymous):

yeah it is an odd question haha @satellite73

OpenStudy (here_to_help15):

No problem :)

Nnesha (nnesha):

ln properties quotient rule\[\large\rm ln y - ln x = \ln \frac{ x }{ y}\] subtraction to division product rule \[\large \rm ln x + \ln y = \ln ( x \times y )\] addition ----> multiplication power rule \[\large\rm ln x^y = y \ln x\] you should apply quotient rule because there is negative sign

OpenStudy (here_to_help15):

Bam @Nnesha you deserve a medal :)

OpenStudy (anonymous):

i was thinking the quotient property but i wasnt too sure :) thanks @Nnesha

OpenStudy (anonymous):

i think there is a mistake in the answer you were given

Nnesha (nnesha):

there is a negative sign so u should be 100000% sure :-)

OpenStudy (anonymous):

what is wrong? @satellite73

OpenStudy (here_to_help15):

Me @satellite73 im anxious where did i make my mistake 0.0

OpenStudy (here_to_help15):

???

OpenStudy (anonymous):

\[3\ln(x)-5\ln(c)\\ \ \ln(x^3)-\ln(c^5)\\ \huge \ln(\frac{x^3}{c^5})\]

OpenStudy (anonymous):

that is assuming it is the question that i wrote above, not something else

OpenStudy (anonymous):

yeah you got it

OpenStudy (anonymous):

that is why i wrote \[\huge 3\ln(x)-5\ln(c)\]first to make sure it was right

OpenStudy (anonymous):

yes :)